OpenStudy (anonymous):

-5/6 + x = -1/2 - 2/3

OpenStudy (anonymous):

\[x = \frac{-1}{2} - \frac{2}{3} + \frac{5}{6} \implies x = \frac{-3 - 4 + 5}{6} \implies \color{green}{x = \frac{-2}{6}} \implies \color{blue}{x = \frac{-1}{3}}\]

hero (hero):

Do you have any idea how he got that result @monica1120 ?

OpenStudy (anonymous):

If not then don;t worry he will explain you far better than me.. After all I am his follower..

OpenStudy (anonymous):

@monica1120 you are allowed to reply.. Do you get what I did above.. And if not then let me know I will use some other method to make you understand..